Consider 0.333 mL of an ideal gas initially at 44.00 °C in a sealed container. What is the final volume after the gas is heated to 100.00 °C at constant pressure?
Answer #1

According to Charle's Law,

V1/T1 = V2/T2

V1 & V2 = Volume

T1 & T2 = absolute temperature

V1 = 0.333 ml. T1 = 44°C. = 273 K + 44. = 317 K

V2 = ? T2 = 100°C. = 273K + 100. = 373 K

0.333/317 = V2/373

0.001051 = V2/373

V2 = 0.001051 × 373

V2 = 0.392 ml

Final volume = 0.392 ml